OpenStudy (anonymous):

Find the slope of the line through the pair of points. A(2, –3), P(2, 9)

OpenStudy (anonymous):

Slope = rise/ run = (y2-y1)/(x2-x1) (9-(-3))/(2-2) = 0 cannot divide by zero. Both points have the same x coordinate which means a vertical line

OpenStudy (anonymous):

so its undefined?

OpenStudy (anonymous):

The slope is 0, it is a straight vertical line.